OpenStudy (anonymous):

can i have some help please, will favourite fan and testimonial, you name it!

OpenStudy (anonymous):

it's question 5

OpenStudy (anonymous):

@ganeshie8 pleaseee

ganeshie8 (ganeshie8):

what do you know about radius of the circle that just touches the x axis ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

radius equals "y" coordinate of the center right ?

OpenStudy (anonymous):

these are my attempts @ganeshie8 but i have a problem at the end there

ganeshie8 (ganeshie8):

paer i looks good!

OpenStudy (anonymous):

what about paper 2? where am i going wrong?

ganeshie8 (ganeshie8):

sry i meant to say "part i" looks good

OpenStudy (anonymous):

oh what about part 2??

ganeshie8 (ganeshie8):

its wrong

ganeshie8 (ganeshie8):

looks you have interpreted the problem incorrectly

OpenStudy (jhannybean):

\[\sf x^2 + 2gx +y^2 +2fy +c=0\]\[\sf (x^2 +2gx+g^2)+(y^2 +2fy +f^2)+c=0\]\[\sf (x+g)^2 +(y+f)^2 + -g^2-f^2+c=0\] Something like that for part i using completing the squares

OpenStudy (anonymous):

oh, how should i have interpreted it?

OpenStudy (anonymous):

ok thnks jhannybean

ganeshie8 (ganeshie8):

you're assuming the given points A and B are centers which is wrong

OpenStudy (anonymous):

i didn't even know i was doing that? how was i even doing that, i just thought i had to find equations to find out the value of the 3 variables g,f, c ??

ganeshie8 (ganeshie8):

R is not |-4| or |-2|

ganeshie8 (ganeshie8):

you need to solve below system : \[3^2+4^2+2g(3)+2f(4) +c = 0 \tag{1}\] \[5^2+2^2+2g(5)+2f(2) +c = 0 \tag{2}\] \[g^2 = c \tag{3}\]

OpenStudy (anonymous):

so do you just sub g^2 into your first 2 equations?

ganeshie8 (ganeshie8):

yes just replace "c" by g^2 in first two equations and solve g, f first

OpenStudy (anonymous):

wait, i sub g^2 into both equations then take the second equation from the 1st one?

ganeshie8 (ganeshie8):

hey no wait, thats really a bad idea

OpenStudy (anonymous):

oh so what should i do then ?

ganeshie8 (ganeshie8):

i can't think of any other easy way, go ahead... do it :)

OpenStudy (anonymous):

yeah i got -4g+4f=0 when i solved them, like earlier on... now what do i do? :)

OpenStudy (jhannybean):

If you're still interested in the completing the square method.. continue from here :) \[\sf (x+g)^2 +(y+f)^2 + -g^2-f^2+c=0\] You've gotten to complete the square but now you're stil wondering, "how do g, f , and c relate to a circle?" We know \(\sf (x+h)^2 +(y+k)^2 =r^2\)- equation of a circle.

OpenStudy (jhannybean):

But in our question, we have \(\sf (x+g)^2 +(y+f)^2 \color{red}{-g^2-f^2+c}=0\) Relating the equation of the circle to our equation, we have:\[\color{blue}{(x+h)^2} +\color{green}{(y+k)^2} =\color{purple}{r^2}\]\[\color{blue}{(x+g)^2}+\color{green}{(y+f)^2}= \color{purple}{g^2+f^2-c}\]

OpenStudy (anonymous):

i'm so confused ....

OpenStudy (anonymous):

@ganeshie8 what do i do after taking them away?

ganeshie8 (ganeshie8):

plugging c = g^2 in first two equations gives u : 25 + 6g + 8f + g^2 = 0 29 + 10g + 4f + g^2 = 0 subtracting both gives you : 4 + 4g - 4f = 0 g = f - 1